OpenStudy (anonymous):

(1, 0) is an ordered pair of the function f(x) = 1 + x.

OpenStudy (hauntedwoodsgal):

Example: Find f(0), f(1), f( - 2), f( - x), - f(x), f(x+1), f(2x), f(x+h) for f(x) = 3x 2 - 2x – 4 f( 0 ) = 3( 0 ) 2 - 2( 0 ) – 4 = - 4 f( 1 ) = 3( 1 ) 2 – 2( 1 ) – 4 = - 3 f( - 2 ) = 3( - 2 ) 2 - 2( - 2 ) - 4= 12 f( - x ) = 3( - x ) 2 - 2( - x ) - 4 = 3x 2 + 2x – 4 f( x+1 ) = 3( x+1 ) 2 - 2( x+1 ) – 4 = 3(x 2 + 2x +1) - 2x - 2 - 4 = 3x 2 + 6x + 3 - 2x – 6= 3x 2 + 4x – 3 f( 2x ) = 3( 2x ) 2 - 2( 2x ) – 4 = 3(4x 2 ) – 4x – 4= 12x 2 – 4x – 4 f( x+h ) = 3( x+h ) 2 - 2( x+h ) – 4 = 3(x 2 + 2xh +h 2 ) – 2x - 2h - 4 = 3x 2 + 6xh + 3h 2 - 2x – 2h – 4 - f(x) = - (3x 2 – 2x – 4) = - 3x 2 + 2x+4

OpenStudy (hauntedwoodsgal):

please give me a medel

OpenStudy (anonymous):

i just did

OpenStudy (hauntedwoodsgal):

thx

OpenStudy (anonymous):

welcome (:

OpenStudy (hauntedwoodsgal):